OpenStudy (anonymous):

Find the domain of the ration function: f(y)=y+7/y^2-7y

OpenStudy (anonymous):

Interval notation btw

OpenStudy (lyrae):

Solve y^2-7y=0 the domain are every number except the answer to the that equation.

OpenStudy (anonymous):

so (-infinity,0)U(0,7)(7, infinity)?

OpenStudy (anonymous):

theres a U after (0,7)

OpenStudy (lyrae):

\[y^2-7y=0 \]\[y(y-7) = 0\]y=0,7 The domain is every number exept 0 and 7 \[-\infty < y < 0 ,\]\[ 0 < y < 7,\]\[7 < y < \infty \]

OpenStudy (lyrae):

So yes :)